OpenStudy (anonymous):

which of the following is the solution to the equation (sqrt5)^8n=125^n+5

OpenStudy (radar):

\[(5^{1/2})^{8n}=125^{n+5}\] have I copied it correctly or is the last term\[125^{n}+5\]

OpenStudy (anonymous):

its the first one but is is square root the 5

OpenStudy (radar):

Yes \[\sqrt{5}=5^{1/2}\] same thing so it becomes:\[5^{4n}=125^{n+5}\]

OpenStudy (radar):

\[5^{4n}=(5^{3})^{n+5}\]\[5^{4n}=5^{3n+15}\]\[4n=3n+15\]n=15

OpenStudy (anonymous):

thank you